# Enhancing Data-Limited Graph Neural Networks by Actively Distilling Knowledge from Large Language Models

### v1

@@ 3.2. Understanding LLM’s Capability
Figure 1 shows the results of the preliminary experiments. From the figure, we can find that the performance of LLMs can be affected by the degree and the homophily ratio of a node. When we use the nodes with more indegrees or higher homophily ratio, the classification accuracy of LLMs is increased significantly. Compared with degrees, LLMs are more sensitive to the homophily ratio. From the right figure of Figure 1, it is obvious that the performance of LLMs changed drastically across nodes with different homophily ratios. For example, the accuracy of LLMs for nodes with the lowest homophily ratio in the Cora dataset is around $40\%$, but it reaches $75\%$ on nodes with the highest homophily ratio.
@@
Our preliminary experiments demonstrate that LLMs are capable of generating high-quality pseudo-labels for nodes with a higher homophily ratio and larger degree, which paves us a way to effectively select nodes to query LLMs to obtain high-quality knowledge for enhancing GNNs knowledge.
Figure 2. An illustration of the proposed framework
3.3. Problem Statement
@@
As LLMs could not give reliable knowledge for all the nodes, in this paper, we study a novel problem of how to effectively leverage LLMs to enhance the performance of few-shot node classification over graphs. Given a text-attributed graph $\mathcal{G}_{T}=(\mathcal{V},\mathbf{A},\mathcal{X})$ with a very limited labeled node set $\mathcal{V}_{l}$ (i.e. $|\mathbf{V}_{l}|\ll|\mathbf{V}_{u}|$) and their label set $\mathcal{Y}_{l}$, a budget size $B$ (note that the budget size $B$ is the number of nodes per class), and a large language model $LLM$, we aim to train a GNN model that can have better performance with only few available labeled nodes by querying $LLM$ within the budget size $B$. The details will be introduced in Section 4.
4. Proposed Model
@@
Though GNNs have shown great power in node classification, the vanilla GNNs suffer from low generalizability with few labeled data for training (Feng et al., 2020). Thus, in order to enhance the generalizability of GNNs, we propose a framework that integrates GNNs with LLMs and employs a novel Graph-LLM-based active learning strategy to actively distill knowledge from LLMs. Our proposed model uses GNN as the backbone model and takes advantage of LLMs’ zero-shot pseudo-labeling and reasoning capabilities, especially for nodes that are difficult for GNN to give accurate predictions. In these instances, LLMs can offer reliable pseudo-labels and provide enhanced rationales, thereby improving the few-shot learning capability of GNNs from distinct perspectives.
@@
An illustration of the framework is shown in Figure 2. Specifically, LLM serves as a teacher model, instructing the student model (GNNs) from two distinct perspectives: (1) it imparts the “correct” answers to the student model along with the probability distribution for all potential categories, drawing upon its vast knowledge, which will teach GNNs with the output logits; and (2) it explains the rationale behind its decision-making process, providing insights into why certain decisions are made, which will serve as the feature teacher to teach GNNs at embedding level. Then the knowledge obtained from LLMs will be distilled to GNNs, and GNNs propagate label information to all unlabeled nodes. We leverage Graph-LLM-based active learning to identify nodes that GNNs struggle to generate reliable pseudo labels but LLMs can provide reliable predictions. These selected nodes are then added to the train set with pseudo labels, and fed to LLMs for logits and rationales, which can further enhance the capability of GNNs under the guidance of LLMs. Finally, we train the ultimate student model, enhancing its generalizability with limited available data. Next, we introduce each component in detail.
4.1. Base GNN Classifier
@@
As GNNs have shown great power in semi-supervised node classification, we adopt Graph Neural Networks (GNNs) as the backbone models, which can be used to capture the structure information between entities and naturally propagate the information to all unlabeled nodes efficiently. We first use SBERT (Reimers and Gurevych, 2019) to encode raw texts $\mathcal{X}$ to text embeddings $\mathbf{X}$. Then, we use GNNs to perform on the given graph and these embeddings. Specifically, the GNN takes the graph $\mathcal{G}_{T}$ as input and learns the node representation as
(1) $\mathbf{H}^{\mathtt{f}}=GNN(\mathbf{A},\mathbf{X})$
@@
(2) $\mathbf{Z}=\text{softmax}(\mathbf{H}^{\mathtt{f}})$
@@
where $\mathbf{Z}$ is the probabilities for all nodes in the graph. The loss function for training the GNN will be introduced in 4.4.
4.2. Obtaining Knowledge from LLM
@@
Despite GNNs showing success in dealing with graph data, the generalizability of GNNs with few available data is still limited. To tackle this challenge, we introduce LLMs as teacher models, leveraging their zero-shot ability (Kojima et al., 2022) to instruct GNNs in classification tasks and provide insights into the reasoning behind these decisions. In this way, GNNs can learn hidden label distribution information and enhanced feature information from LLMs, which empower the capabilities of GNNs with scarce labeled data
@@
To effectively distill knowledge from LLMs to GNN, we consider two types of knowledge: (1) soft labels and logits; and (2) rationales behind LLMs’ decision-making process. Soft labels and logits reveal hidden distribution information for unlabeled data, while rationales contribute richer node information. This combination allows GNNs to benefit from the unlabeled data and get enhanced node features. We prompt the prediction and reasoning in a two-step manner: first, we input the raw texts into the LLMs to generate the soft labels and logits with the probability distribution. We then let LLMs explain the reason why they make these decisions. Examples of prompts are shown in Table 1. Due to the possibility of LLMs producing outputs that deviate from the desired format when prompts contain numerous tasks, thereby increasing uncertainty in the parsing process, we propose to utilize two separate prompts to obtain logits and rationales. Next, we give the details.
Table 1. Prompt examples
@@
For soft labels and logits Paper: ¡Paper Information¿. Task: There are following categories: ¡categories¿, which categories does this paper belong to? Provide your ¡k¿ best guesses within the given categories: ¡categories¿ and a confidence score that each is correct (0 to 1). The sum of all confidence should be 1. Outputs must be in the given categories. For example: ”answer”: ¡your first answer¿, ”confidence”: ¡confidence for first answer¿ , …
@@
For rationales Paper: ¡Paper Information¿. Task: There are following categories: ¡categories¿, which categories does this paper belong to? Let’s think step by step. Explain your decision in detail.
4.2.1. Soft Labels and Logits Generation
@@
For a node $v_{i}$, our proposed model first feeds raw text $\mathcal{X}_{i}$ into LLMs to generate soft labels $\bar{y}_{i}$ for $\mathcal{X}_{i}$ and the logits $\mathbf{l}_{i}$ for all possible categories. An example of the prompt for soft labels and logit generation is shown in the first row of Table 1. We leverage the zero-shot ability of LLMs to generate relevant reliable soft labels and logits so that GNNs can leverage the hidden information of unlabeled data with knowledge distillation. This can be formally written as
(3) $\bar{y}_{i},\mathbf{l}_{i}=LLM(\mathcal{X}_{i};prompt)$
4.2.2. Rationales for Feature Enhancement
@@
Traditional knowledge distillation methods primarily utilize the soft labels and logits from the teacher model. Nonetheless, incorporating the rationales behind text decisions can significantly enhance the learning capabilities of GNNs (He et al., 2023). In this context, GNNs are able to learn more informative features from the LLM at the embedding level. Consequently, we introduce LLMs as a feature teacher, guiding GNNs to assimilate more informative features in their decision-making process. Different from previous works that concatenate the enhanced embeddings and the node embeddings or simply replace the node representations directly, we will use a loss function to minimize the difference between them, which will help GNNs learn the enhanced representation while retaining the original node representations. The loss function will be detailed in Section 4.3
@@
(4) $\mathcal{R}_{i}=LLM(\mathcal{X}_{i};prompt)$
An example of the prompt for rationales is shown in the second row of Table 1. Since the rationales we get from LLMs are all textual explanations, we further need to transform them into the embedding level to teach GNNs the more informative features. We use a pre-trained language model such as Sentence BERT (SBERT) (Reimers and Gurevych, 2019) to get the embeddings for $\mathcal{R}_{i}$, which can be represented as follows:
(5) $\bar{r}_{i}=LM(\mathcal{R}_{i})$
where $\bar{r}_{i}$ means the embedding of $i$-th rationale.
However, since the dimensionality of $\bar{r}_{i}$ may differ from the dimension of the final layer of GNNs, alignment between these representations is necessary. While min/max pooling can effectively reduce dimensionality for alignment purposes, it tends to lose information during the pooling process. To retain the enriched information from these rationales, we train a Multi-Layer Perceptron (MLP) using text embeddings $\mathbf{X}_{l}$ of the limited labeled node set $\mathcal{V}_{l}$ and their corresponding ground truth labels $\mathcal{Y}_{l}$, applying the cross-entropy loss function. This MLP is tasked with aligning the representations between the rationales $\bar{r}$ and the outputs $\mathbf{H}^{\mathtt{f}}$ of the final layer of GNNs, ensuring that valuable information is retained throughout the alignment process. The final representation for $i$-th rationale is generated as follows:
(6) $r_{i}=MLP(\bar{r}_{i})$
@@
4.3. Distilling Knowledge to GNN
With the knowledge from LLM represented as $\mathbf{r}_{i}$ and $\mathbf{l}_{i}$, we use knowledge distillation (Hinton et al., 2015) to distill this knowledge into GNNs. Through this process, GNNs can tap into the hidden information behind unlabeled nodes by using output logits to enhance their performance. Moreover, they can achieve improved node representations by incorporating the rationales generated from LLMs to further enrich the depth and quality of the information being processed. Specifically, LLMs serve as a pre-trained teacher model to teach the student model (GNNs) from two distinct perspectives: 1). soft labels and the probability distribution (logits) and 2). the rationales at the embedding level.
4.3.1. Loss for Knowledge Distillation
Let $\mathcal{V}_{S}$ be the set of nodes including the original training data and the data selected through active learning (to be introduced in Section 4.5). Following (Li et al., 2022), for each $v_{i}\in\mathcal{V}_{s}$, we first convert the logits $\mathbf{l}_{i}$ from LLM as:
(7) $p(y_{i}=j|LLM)=\frac{\exp{\left(\mathbf{l}_{ij}/{\tau}\right)}}{\sum_{c=1}^{C}\exp{\left(\mathbf{l}_{ic}/{\tau}\right)}}$
where $C$ is the number of classes, $\tau$ is the knowledge distillation (KD) temperature to control how much of the teacher’s knowledge is distilled to student model and $\mathbf{l}_{ij}$ is the $j$-th elements of $\mathbf{l}_{i}$. Then, the student can learn the distilled knowledge from the teacher by optimizing the following loss function:
(8) ${\mathcal{L}}_{T}=-\frac{1}{|{\mathcal{V}}_{S}|}\sum_{v_{i}\in{\mathcal{V}}_{S}}\sum_{j=1}^{C}p(y_{i}=j|LLM)\log Z_{ij}$
@@ 4.3.1. Loss for Knowledge Distillation
4.3.2. Loss for Feature Alignment
We also introduce rationales to augment the node representation from a feature perspective at the embedding level. With the embeddings $\mathbf{r}$ we get from 4.2, Mean Square Error (MSE) is used to calculate the loss between the rationales and node embeddings $\mathbf{H}^{\mathtt{f}}$ at the final layer of GNNs for all the nodes in the current training set ${\mathcal{V}}_{S}$ as:
(9) ${\mathcal{L}}_{F}=\frac{1}{|{\mathcal{V}}_{S}|}\sum_{i=1}^{|{\mathcal{V}}_{S}|}(\mathbf{H}^{\mathtt{f}}_{i}-\mathbf{r}_{i})^{2}$
By employing this approach, the GNNs learn informative rationales from LLMs, enhancing their learning capabilities from a feature perspective at the embedding level.
4.4. Objective Function of Proposed Framework
The student model itself computes training loss between predictions and ground truth (hard attribute label), which is defined as:
(10) ${\mathcal{L}}_{S}=-\frac{1}{|{\mathcal{V}}_{S}|}\sum_{v_{i}\in\mathcal{V}_{S}}\sum_{c=1}^{C}\mathbb{I}(y_{i}=c)\log Z_{ic}$
@@ 4.4. Objective Function of Proposed Framework
(11) $\min_{GNN}{\mathcal{L}}=(1-\alpha-\beta){\mathcal{L}}_{S}+\alpha{\mathcal{L}}_{T}+\beta{\mathcal{L}}_{F}$
@@
4.5. Graph-LLM-based Active learning

### v2

@@ III-A Understanding LLM’s Capability
Figure 1 shows the preliminary experimental results. From the figure, we can find that the performance of LLMs can be affected by the degree and the homophily ratio of a node. When we use the nodes with more indegrees or higher homophily ratio, the classification accuracy of LLMs is increased significantly. Compared with degrees, LLMs are more sensitive to the homophily ratio. From the right figure of Figure 1, it is obvious that the performance of LLMs changed drastically across nodes with different homophily ratios. For example, the accuracy of LLMs for nodes with the lowest homophily ratio in the Cora dataset is around $40\%$, but it reaches $75\%$ on nodes with the highest homophily ratio. This is because nodes with more degrees and a higher homophily ratio tend to be closer to the distribution center and occupy well-connected positions in the graph. These nodes exert indirect influence and are naturally associated with richer textual information, making them more significant and representative within clusters. For instance, in citation networks, papers with more citations are more representative and easily distinguishable within their field. These representative and distinct contexts lead LLMs to achieve better classification outcomes.
@@
Our preliminary experiments demonstrate that LLMs are capable of generating high-quality pseudo-labels for nodes with a higher homophily ratio and more degree, which paves us a way to effectively select nodes to query LLMs to obtain high-quality knowledge for enhancing GNNs knowledge.
Figure 2: An illustration of the proposed framework
III-B Problem Statement
@@
As LLMs could not give reliable knowledge for all the nodes, in this paper, we study a novel problem of how to effectively leverage LLMs to enhance the performance of few-shot node classification over graphs. Given a text-attributed graph $\mathcal{G}_{T}=(\mathcal{V},\mathbf{A},\mathcal{X})$ with a very limited labeled node set $\mathcal{V}_{l}$ (i.e. $|\mathcal{V}_{l}|\ll|\mathcal{V}_{u}|$) and their label set $\mathcal{Y}_{l}$, a budget size $B$ (note that the budget size $B$ is the number of nodes per class), and a large language model $LLM$, we aim to train a GNN that can have better performance with only few available labeled nodes by querying $LLM$ within the budget $B$.
IV Proposed Model
@@
Though GNNs have shown great power in node classification, the vanilla GNNs suffer from low generalizability with few labeled data for training [41]. Thus, in order to enhance the generalizability of GNNs, we propose a framework that integrates GNNs with LLMs and employs a novel Graph-LLM-based active learning strategy to actively distill knowledge from LLMs. Our proposed model uses GNN as the backbone model and takes advantage of LLMs’ zero-shot pseudo-labeling and reasoning capabilities, especially for nodes that are difficult for GNN to give accurate predictions. In these instances, LLMs can offer reliable pseudo-labels and provide enhanced rationales, thereby improving the few-shot learning capability of GNNs from distinct perspectives.
@@
An illustration of the framework is shown in Figure 2. Specifically, LLM serves as a teacher model, instructing the student model (GNNs) from two distinct perspectives: (1) it imparts the “correct” answers to the student model along with the probability distribution for all potential categories, drawing upon its vast knowledge, which teaches GNNs with the output logits; and (2) it explains the rationale behind its decision-making process, providing insights into why certain decisions are made, which serves as the feature teacher to teach GNNs at embedding level. Then the knowledge obtained from LLMs will be distilled to GNNs, and GNNs propagate label information to all unlabeled nodes. We leverage Graph-LLM-based active learning to identify nodes that GNNs struggle to generate reliable pseudo labels but LLMs can provide reliable predictions. These selected nodes are then added to the train set with pseudo labels, and fed to LLMs for logits and rationales, which can further enhance the capability of GNNs under the guidance of LLMs. Finally, we train the ultimate student model, enhancing its generalizability with limited data. Next, we introduce each component in detail.
IV-A Base GNN Classifier
@@
As GNNs have shown great power in semi-supervised node classification, we adopt Graph Neural Networks (GNNs) as the backbone models, which can be used to capture the structure information between entities and naturally propagate the information to all unlabeled nodes efficiently. We first use SBERT [42] to encode raw texts $\mathcal{X}$ to text embeddings $\mathbf{X}$. Then, we use GNNs to perform on the given graph and these embeddings. Specifically, the GNN takes the graph $\mathcal{G}_{T}$ as input and learns the node representation as
$\mathbf{H}^{\mathtt{f}}=GNN(\mathbf{A},\mathbf{X})$ (1)
@@
$\mathbf{Z}=\text{softmax}(\mathbf{H}^{\mathtt{f}})$ (2)
@@
where $\mathbf{Z}$ is the probabilities for all nodes in the graph. The loss function for training the GNN will be introduced in IV-D.
IV-B Obtaining Knowledge from LLM
@@
Despite GNNs showing success in dealing with graph data, the generalizability of GNNs with few available data is still limited. To tackle this challenge, we introduce LLMs as teacher models, leveraging their zero-shot ability [20] to instruct GNNs in classification tasks and provide insights into the reasoning behind these decisions. In this way, GNNs can learn hidden label distribution information and enhanced feature information from LLMs, which empower the capabilities of GNNs with scarce labeled data
@@
To effectively distill knowledge from LLMs to GNN, we consider two types of knowledge: (1) soft labels and logits; and (2) rationales behind LLMs’ decision-making process. Soft labels and logits reveal hidden distribution information for unlabeled data, while rationales contribute richer node information. This combination allows GNNs to benefit from the unlabeled data and get enhanced node features. We prompt the prediction and reasoning in a two-step manner: first, we input the raw texts into the LLMs to generate the soft labels and logits with the probability distribution. We then let LLMs explain the reason why they make these decisions. Examples of prompts are shown in Table I. To avoid deviations in output formats with multi-task prompts, we use separate prompts for logits and rationales, respectively. Next, we give the details.
TABLE I: Prompt examples
@@
For soft labels and logits Paper: $<$Paper Information$>$. Task: For the following categories: $<$categories$>$, which categories does this paper belong to? Provide your $<$k$>$ best guesses within the given categories: $<$categories$>$ and a confidence score that each is correct (0 to 1). The sum of all confidence should be 1. Outputs must be in the given categories. For example: “answer”: $<$your first answer$>$, “confidence”: $<$confidence for first answer$>$, …
@@
For rationales Paper: $<$Paper Information$>$. Task: For the following categories: $<$categories$>$, which categories does this paper belong to? Think step by step. Explain your decision in detail.
IV-B1 Soft Labels and Logits Generation
@@
For a node $v_{i}$, we first feeds raw text $\mathcal{X}_{i}$ into LLMs to generate soft labels $\bar{y}_{i}$ for $\mathcal{X}_{i}$ and the logits $\mathbf{l}_{i}$ for all possible categories. An example of the prompt for soft labels and logit generation is shown in the first row of Table I. We leverage the zero-shot ability of LLMs to generate relevant reliable soft labels and logits so that GNNs can leverage the hidden information of unlabeled data with knowledge distillation. This can be written as
$\bar{y}_{i},\mathbf{l}_{i}=LLM(\mathcal{X}_{i};prompt)$ (3)
IV-B2 Rationales for Feature Enhancement
@@
Traditional knowledge distillation methods primarily utilize the soft labels and logits from the teacher model. Nonetheless, incorporating the rationales behind text decisions can significantly enhance the learning capabilities of GNNs [22]. In this context, GNNs are able to learn more informative features from the LLM at the embedding level. Consequently, we introduce LLMs as a feature teacher, guiding GNNs to assimilate more informative features in their decision-making process. Unlike previous works that concatenate the enhanced embeddings and the node embeddings or simply replace the node representations directly, we will use a loss function to minimize the difference between them, which will help GNNs learn the enhanced representation while retaining the original node representations. The loss function will be detailed in Section IV-C.
@@
$\mathcal{R}_{i}=LLM(\mathcal{X}_{i};prompt)$ (4)
An example of the prompt for rationales is shown in the second row of Table I. Since the rationales we get from LLMs are all textual explanations, we further need to transform them into the embedding level to teach GNNs the more informative features. We use a pre-trained language model such as Sentence BERT (SBERT) [42] to get the embeddings for $\mathcal{R}_{i}$, which can be represented as follows:
$\bar{\mathbf{r}}_{i}=LM(\mathcal{R}_{i})$ (5)
where $\bar{\mathbf{r}}_{i}$ means the embedding of $i$-th rationale.
However, since the dimensionality of $\bar{\mathbf{r}}_{i}$ may be different from the dimension of the final layer of GNNs, alignment between these representations is necessary. While min/max pooling can effectively reduce dimensionality for alignment purposes, it tends to lose information during the pooling process. To retain the enriched information from these rationales, we train a Multi-Layer Perceptron (MLP) using text embeddings $\mathbf{X}_{l}$ of the limited labeled node set $\mathcal{V}_{l}$ and their corresponding ground truth labels $\mathcal{Y}_{l}$, applying the cross-entropy loss function. This MLP is tasked with aligning the representations between the rationales $\bar{\mathbf{r}}$ and the outputs $\mathbf{H}^{\mathtt{f}}$ of the final layer of GNNs, ensuring that valuable information is retained throughout the alignment process. The final representation for $i$-th rationale is generated as follows:
$\mathbf{r}_{i}=MLP(\bar{r}_{i})$ (6)
@@
IV-C Distilling Knowledge to GNN
With the knowledge from LLM represented as $\mathbf{r}_{i}$ and $\mathbf{l}_{i}$, we use knowledge distillation [43] to distill this knowledge into GNNs. Through this process, GNNs can tap into the hidden information behind unlabeled nodes by using output logits to enhance their performance. Moreover, they can achieve improved node representations by incorporating the rationales generated from LLMs to further enrich the depth and quality of the information being processed. Specifically, LLMs serve as a pre-trained teacher model to teach the student model (GNNs) from two distinct perspectives: 1). soft labels and the probability distribution (logits) and 2). the rationales at the embedding level.
IV-C1 Loss for Knowledge Distillation
Let $\mathcal{V}_{S}$ be the set of nodes including the original training data and the data selected through active learning (to be introduced in Section IV-E). Following [5], for each $v_{i}\in\mathcal{V}_{s}$, we first convert the logits $\mathbf{l}_{i}$ from LLM as:
$p(y_{i}=j|LLM)=\frac{\exp{\left(\mathbf{l}_{ij}/{\tau}\right)}}{\sum_{c=1}^{C}\exp{\left(\mathbf{l}_{ic}/{\tau}\right)}}$ (7)
where $C$ is the number of classes, $\tau$ is the knowledge distillation (KD) temperature to control how much of the teacher’s knowledge is distilled to student model and $\mathbf{l}_{ij}$ is the $j$-th elements of $\mathbf{l}_{i}$. Then, the student can learn the distilled knowledge from the teacher by optimizing the following loss:
${\mathcal{L}}_{T}=-\frac{1}{|{\mathcal{V}}_{S}|}\sum_{v_{i}\in{\mathcal{V}}_{S}}\sum_{j=1}^{C}p(y_{i}=j|LLM)\log Z_{ij}$ (8)
@@ IV-C1 Loss for Knowledge Distillation
IV-C2 Loss for Feature Alignment
We also introduce rationales to augment the node representation from a feature perspective at the embedding level. With $\mathbf{r}$ we get from IV-B, Mean Square Error (MSE) is used to calculate the loss between the rationales and node embeddings $\mathbf{H}^{\mathtt{f}}$ at the final layer of GNNs for all the nodes in the current training set ${\mathcal{V}}_{S}$ as:
${\mathcal{L}}_{F}=\frac{1}{|{\mathcal{V}}_{S}|}\sum_{i=1}^{|{\mathcal{V}}_{S}|}(\mathbf{H}^{\mathtt{f}}_{i}-\mathbf{r}_{i})^{2}$ (9)
where $\mathbf{H}^{\mathtt{f}}_{i}$ is the node embedding of $v_{i}$ from GNN. By employing this approach, the GNNs learn informative rationales from LLMs, enhancing their learning capabilities from a feature perspective at the embedding level.
IV-D Objective Function of Proposed Framework
The student model itself computes training loss between predictions and labeled or pseudo-labeled data as
${\mathcal{L}}_{S}=-\frac{1}{|{\mathcal{V}}_{S}|}\sum_{v_{i}\in\mathcal{V}_{S}}\sum_{c=1}^{C}\mathbb{I}(y_{i}=c)\log Z_{ic}$ (10)
@@ IV-D Objective Function of Proposed Framework
$\min_{GNN}{\mathcal{L}}=(1-\alpha-\beta){\mathcal{L}}_{S}+\alpha{\mathcal{L}}_{T}+\beta{\mathcal{L}}_{F}$ (11)
@@
IV-E Graph-LLM-based Active learning